Suppose that you have just borrowed $250,000 in the form of a 30 year mortgage. The loan has an annual interest rate of 9% wineed help thanks!
Answer #1
Question1
Pv Amount of Loan $250,000
Rate Monthly interest =(9/12)% 0.75%
Nper Number of months=30*12 360
a PMT Monthly Payment for the loan $2,011.56 (Using PMT function of excel with rate=0.75%, Nper=360, Pv=-250000)
b Loan Balance at end of year 12=Present Value of monthly payments during year13-30(18*12=216 months)
Rate Monthly interest =(9/12)% 0.75%
Nper Number of months=18*12 216
PMT Monthly Payment for the loan $2,011.56
PV Loan Balance at end of year 12 $214,807.80 (Using PV function of excel with rate=0.75%, Nper=216, Pmt=-2011.56)
c Interest You Pay in 7th Year
Loan Balance at end of year 6 $237,026.17 (Using PV function of excel with rate=0.75%, Nper=(360-72), Pmt=-2011.56)
Loan Balance at end of year 7 $234,101.14 (Using PV function of excel with rate=0.75%, Nper=(360-84), Pmt=-2011.56)
Total Principal Payment in Year 7 $2,925.03 (237026.17-234101.14)
Total amount of payment during 7th year $24,138.68 (2011.56*12)
Total Interest Payment during 7th year $21,213.65 (24138.68-2925.03)
d 248th payment principal
Loan Balance at end of 247th payment $152,920.69 (Using PV function of excel with rate=0.75%, Nper=(360-247), Pmt=-2011.56)
Interest Component of 248th payment $1,146.91 (0.75%*152920.69)
Monthly Payment $2,011.56
Principal Component of 248th Payment $864.65 (2011.56-1146.91)
=PMT(F5,F6,-F4) F7 A В C D Е F G Н К L M 1 2 Question1 $250,000 Amount of Loan 4 Pv Monthly interest =(9/12)% 0.75% 5 Rate Nufox =PV(F11, F12,- F13) F14 A В D Е F G H к L м 1 2 Question1 $250,000 Amount of Loan 4 Pv Monthly interest =(9/12)% 0.75% 5F17 =PV(F11,288,-F13) A В C E F Н K L 1 2 3 Question1 Amount of Loan $250,000 4 Pv Monthly interest =(9/12)% 0.75% 5 Rate Numfix F18 =PV(F11,276,-F7) A B C D Е F G H K L M 1 2 3 Question1 $250,000 Amount of Loan 4 Pv Monthly interest (9/12)% 0.75% 5=PV (F11,113,-F7) X F24 А В C D Е F G Н J K м 1 2 3 Question1 $250,000 Amount of Loan 4 Pv Monthly interest =(9/12)% 0.75% 5